OpenStudy (anonymous):

solve the following quadratic equations by factoring: 2x^2-3x=x^2+18

OpenStudy (hba):

First of all we need to bring it into the standard form of quadratic equation and compare to find a,b and c The standard form is, \[ax^2+bx+c=0\] Then we need to use the quad formula which is, \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] The other methods we can use here are as follows: 1)Breaking the middle term 2)Completing the whole square.

OpenStudy (anonymous):

but my problem is exactly where to start

OpenStudy (anonymous):

X(X-3) + 6(X-3) = 0

OpenStudy (anonymous):

X=3

OpenStudy (hba):

Um,The first step is to bring it into standard form, \[ax^2+bx+c=0\]

OpenStudy (anonymous):

so we would have............

OpenStudy (hba):

2x^2-3x=x^2+18 Subtract both sides by (x^2+18)

OpenStudy (anonymous):

Sorry is X(X+3) - 6(X+3) = 0 X= +/- 3

OpenStudy (anonymous):

\[x^2-3x-18=0\]

OpenStudy (hba):

Yeah right :)

OpenStudy (anonymous):

so now im stuck again........hmmph

OpenStudy (hba):

Can you use the quadratic formula ?

OpenStudy (anonymous):

we divide by x ???

OpenStudy (hba):

Think of two numbers when we add or subtract we get -3 and when we multiply we get -18

OpenStudy (anonymous):

i know its 6 but what would the work looklike

OpenStudy (hba):

What do you mean by its 6 ?

OpenStudy (hba):

I said two numbers.

OpenStudy (anonymous):

-3 and 6

OpenStudy (hba):

Ok so here it is, \[x^2-6x+3x-18=0\]

OpenStudy (hba):

Now take common from first two terms and then last two terms.

OpenStudy (anonymous):

\[so -6x+3x would be together on 1 term\]

OpenStudy (hba):

No like, x(x-6) + 3 (x-6)=0

OpenStudy (hba):

Either, x-6=0 Or, x+3=0

OpenStudy (hba):

Now solve and get the values of x :)

OpenStudy (anonymous):

thank you so much for the help